# Is Complex Query Answering Really Complex?

### v2

@@ 6 New Benchmarks for Undistorted CQA Performance
Building new CQA benchmarks. We generate our benchmarks to comprise both full-inference and partial-inference QA pairs, but we modify the algorithm of (Ren et al., 2020) to ensure that no training links are present in the reasoning trees of full-inference pairs and we balance each query type a QA pair can be reduced to by sampling 10,000 QA pairs each. For example, for 3p queries, we sample 30.000 QA pairs, 10.000 for 3p that can be reduced to 1p, 10.000 for those reduceable to 2p, and 10.000 for full-inference QA pairs (3p QA pairs non-reduceable to any other type). Moreover, to raise the bar of “complexity” in CQA, we introduce two query types that, in their full-inference versions, are harder than simpler types by design. Specifically, we design “4p” queries for a sequential path made of four links and “4i”, which represents the intersection of the target entity sets defined over four links. For their logical formulation, see App. E
@@
How do SoTA methods fare on our new benchmarks? For $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$ we re-used the pre-trained models used for the experiments in Sec. 4, while for the newly created $\mathsf{ICEWS18}$+$\mathsf{H}$ we trained GNN-QE, ConE, and the link predictor used in CQD, CQD-Hybrid and QTO from scratch. As this is not necessary for ULTRAQ, being a zero-shot neural link prediction applicable to any KG, we re-used the checkpoint provided by the authors. Hyperparameters of all models are in Sec. F.2.
@@
Table 4 shows the results of the selected baselines on the new benchmarks, for all query types in Figure 2. For stratified results on $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$ see Tables D.2, D.3 and D.4, respectively. The surprising result from Table 4 is that no model obtains SoTA performance on most query type, which was not the case for the old benchmarks, where QTO obtained SoTA performance for most query types (see Tables A.6 and A.5). This is further evidence that current SoTA methods mainly focused on improving the performance on the easiest QA pairs and likely overfit the benchmarks. Instead, with our new benchmarks, this will not be possible in the same way, as we balanced the amount of partial- and full-inference QA pairs per type.
@@
Similarly, for queries including negation, the overall MRR greatly drops, and no existing model obtains SoTA for every query type, as shown in Table D.5. Furthermore, we highlight how full-inference 2u queries have similar performance than 1p, similar to how full-inference 2u1p are expected to be as hard as 2p, as discussed in Sec. 3. Finally, we remark that $\mathsf{ICEWS18}$+$\mathsf{H}$ is much harder than $\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{FB15k237}$+$\mathsf{H}$, across all query types, even 1p, thus raising the bar for neural link predictors as well. This confirms our hypothesis that the sampling method of the KG splits plays a big role in determining the hardness of the benchmark.
@@
Table 4: There is no clear SoTA method for the new benchmarks. As shown in Figure 3, the MRR on the new benchmarks is significantly lower than the old ones. For example, for 3i queries on $\mathsf{FB15k237}$+$\mathsf{H}$, QTO has an MRR of 10.1, while for $\mathsf{FB15k237}$ it was 54.6. For the stratified analysis see Tables D.2, D.3 and D.4. For results on queries involving negation see Table D.5.
Model 1p 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p 4p 4i
$\mathsf{FB15k237}$+$\mathsf{H}$ GNN-QE 42.8 5.2 4.0 6.0 8.8 5.6 9.9 32.5 10.0 4.3 20.0
ULTRAQ 40.6 4.5 3.5 5.2 7.2 5.3 10.1 29.4 8.3 3.8 16.4
CQD 46.7 4.4 2.4 11.3 12.8 6.0 11.5 40.1 10.6 1.1 23.8
CQD-Hybrid 46.7 4.8 3.1 6.0 8.6 5.5 12.9 42.2 12.0 2.4 17.4
ConE 41.8 4.6 3.9 9.1 10.3 3.8 7.9 22.8 6.0 3.5 20.3
QTO 46.7 4.9 3.7 8.7 10.1 6.1 13.5 30.6 11.2 3.9 20.2
$\mathsf{NELL995}$+$\mathsf{H}$ GNN-QE 53.6 8.0 6.0 10.7 13.3 16.0 13.5 47.5 9.8 4.7 19.4
ULTRAQ 38.9 6.1 4.1 7.9 10.2 15.8 9.3 28.1 9.5 4.2 15.6
CQD 60.4 9.6 4.2 18.5 19.6 18.9 22.6 46.3 18.5 2.0 24.8
CQD-Hybrid 60.4 9.0 6.1 12.1 14.4 17.4 21.2 46.4 19.3 3.5 20.4
ConE 53.1 7.9 6.7 21.8 23.6 14.9 11.8 39.9 8.8 5.2 27.6
QTO 60.3 9.8 8.0 14.6 15.8 17.6 21.1 49.1 18.9 7.0 20.9
$\mathsf{ICEWS18}$+$\mathsf{H}$ GNN-QE 12.2 0.9 0.5 16.1 26.5 19.1 3.5 7.6 1.1 0.4 34.0
ULTRAQ 6.3 1.2 1.2 7.0 11.7 8.8 1.3 3.3 1.2 0.8 15.9
CQD 16.6 2.5 1.5 13.0 19.5 17.1 6.7 6.8 5.9 1.1 24.0
CQD-Hybrid 16.6 2.6 1.5 15.0 25.5 17.5 5.8 6.8 5.6 0.9 33.2
ConE 3.5 0.9 0.9 1.2 0.5 1.2 1.6 1.1 0.9 0.6 0.3
QTO 16.6 2.6 1.4 15.7 25.0 18.4 6.2 6.7 4.9 1.1 31.5
@@
Takeaway 3. Our benchmarks have a balanced amount of partial-inference and full-inference QA pairs that depends on the query type, allowing to measure CQA performance while not distorting the aggregate performance across all QA pairs with different hardness. Additionally, $\mathsf{ICEWS18}$+$\mathsf{H}$ highlights that more realistic sampling strategies are more challenging for current SoTA.
@@ 7 Conclusion
In this paper, we revisit CQA on KGs and reveal that the “good” performance of SoTA approaches predominantly comes from answers that can be reduced to easier types (Table 2), the vast majority of which boiling down to single link prediction (Table 1). We perform such an analysis by dissecting the arbitrary assumptions that were used to create these datasets and by highlighting how current neural and hybrid solvers can exploit (different) forms of triple memorization to make complex queries much easier. We confirm this by reporting their performance in a stratified analysis and by proposing our hybrid solver, CQD-Hybrid, which, while being a simple extension of an old method like CQD, can be very competitive against other SoTA models. This reveals that $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ are not suitable to precisely assess the capability of CQA methods to answer complex queries.
We then created a set of new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$$\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{ICEWS18}$+$\mathsf{H}$ that comprise a balanced amount of full-inference and partial-inference QA pairs. We consider them to be a stepping stone towards benchmarking truly complex reasoning with ML models. However, both the old and our new benchmarks only consider queries with bindings of single target variables. While this reflects an important class of queries on real-world KGs, many real-world queries require bindings to multiple target variables (i.e. answer tuples). We plan to extend our current study of the “real hardness” of CQA benchmarks to other popular settings in neural query answering, such as inductive scenarios (Galkin et al., 2022) where some entities and/or relations are unseen during the test stage.
@@ Acknowledgements
AV was supported by the “UNREAL: Unified Reasoning Layer for Trustworthy ML” project (EP/Y023838/1) selected by the ERC and funded by UKRI EPSRC. PM was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP. CG was funded by SMARTY (grant no. 101140087). BX was partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-1574 – 471687386. DH was supported by the Deutsche Forschungsgemeinschaft (DFG): Germany’s Excellence Strategy – EXC 2120/1, GA 390831618 (RP20).
@@ A.1 Influence of existing links on queries involving negation
To show that our analysis can be valuable also for different kinds of queries, in Table A.1 we show an analysis for queries involving negation. For such analysis, we split the reasoning tree, defined in Sec. 3, into two subparts, namely positive reasoning tree, composed by the non-negated triples, and the negative reasoning tree, only composed by the negated triples. In particular, in the same spirit of Table 1, by only looking at the positive reasoning tree, we report the percentage of $(q,a)$ pairs that can be reduced to easier types (partial-inference) and the one that cannot be reduced to a simpler type (full-inference). Furthermore, we argue that also the negative reasoning tree contains training triples, but how this propagates to performance is less clear than the positive part, as each method treats negation differently. Table A.1 shows that in both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, queries can be reduced to easier types, revealing that our analysis also extends to negated queries. The only exceptions are 2in, 2nu1p where the positive reasoning tree consists of a single link, see App. B. Consequently, no reduction to partial inference is possible (see Sec. 3).
@@
Refer to captionRefer to caption
Figure A.1: Some answers of union queries can be reached by a single missing link, while the other link does not exist as shown for this 2u query and fragments of the KG $\mathsf{FB15k237}$, where $r_{1}=\mathsf{releasedIn}$, $r_{2}=\mathsf{crewMemberOf}$. (Top) Example answer reachable by two missing links. (Bottom) Example answers only reachable by one missing link.
Table A.3: Exploiting the links in the training KG during inference boosts MRR on existing benchmarks, as shown for our CQD-Hybrid model compared to previous SoTA, for several query types. CQD-Hybrid always achieves higher MRR scores when compared to CQD, and outperforms more sophisticated baselines 10/14 times. Best values in bold and second best underlined. For the stratified comparison of CQD-Hybrid with previous SoTA, please refer to the Tables A.5 and A.6.
Method $\mathsf{FB15k237}$ $\mathsf{NELL995}$
2p 3p 2i 3i 1p2i 2i1p 2u1p 2p 3p 2i 3i 1p2i 2i1p 2u1p
GNN-QE 14.7 11.8 38.3 54.1 31.1 18.9 9.7 17.9 15.2 40.0 50.9 29.1 20.5 8.8
ULTRAQ 11.5 8.9 35.7 51.0 29.6 18.6 7.3 11.2 9.7 36.3 47.7 25.1 15.6 8.4
CQD 13.2 7.8 35.0 48.5 27.5 20.7 10.5 22.0 13.4 42.2 51.8 31.5 25.8 16.0
ConE 12.8 11.0 32.6 47.3 25.5 14.0 7.4 16.0 13.8 39.6 50.2 26.1 17.6 11.1
CQD-Hybrid (ours) 15.0 11.0 37.6 52.7 31.2 24.0 10.3 23.8 17.8 44.2 57.8 33.2 28.4 16.6
@@ A.2 Additional comparisons
Table A.6 and Table A.5 show the full results of $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, including the results comparing all methods, all query types, and the subtypes they are reduced to. Furthermore, in Table A.3 we show that our baseline, CQD-hybrid, achieves SoTA performance when compared to non-hybrid baselines, providing evidence that memorization of existing triples leads to inflated performance in the old benchmarks.
Moreover, A.4 show the full results of the union queries, including the old overall results, which include $(q,t)$ pairs having non-existing links in their reasoning tree (see Figure A.1), and the “new overall”, in which those $(q,t)$ pairs are filtered out. In particular, for “2u1p” queries, we observe that for $\mathsf{FB15k237}$, the full-inference queries have a higher MRR than the overall, while for $\mathsf{NELL995}$, we find the MRR results on full-inference $(q,t)$ pairs to be lower than the new overall results as expected. We attribute the results on $\mathsf{FB15k237}$ to the presence of bias in the $(q,t)$ pairs due to their scarcity in the benchmarks. We argue that for the same reason, 2u1p performances are not comparable to 2p. However, we will show that for the full-inference $(q,t)$ pairs in the new benchmark, our claim is confirmed (see Tables D.2, D.3 and D.4).
@@
Table A.4: Full inference 2u $(q,t)$ pairs show higher performance than non-filtered, while full-inference 2u1p pairs show lower or comparable performance. Results on union queries for existing benchmarks and comparison between old and new overalls. ‘-’ refers to reductions that are not possible, while ‘/’ to reductions that are possible but that are not present in the data.
@@
Table A.5: The best model on all available queries (column “overall”) and the best on the full inference queries (diagonal) is different for 6/7 query types, excluding 1p and 2u, being already of type full-inference. Comparison of MRR scores for SoTA methods for the different subtypes $\mathsf{NELL995}$ queries can be reduced to. Best in bold.
@@
Table A.6: The best model on all available queries (column “overall”) and the best on the full inference queries (diagonal) is different for 5/7 query types, excluding 1p and 2u, those (q,a) pairs being all of type full-inference. Comparison of MRR scores for SoTA methods for the different subtypes $\mathsf{FB15k237}$ queries can be reduced to. Best in bold.
@@
The stratified MRR performance of QTO (Bai et al., 2023) for the old benchmarks have been included in Tables A.6 and A.5. For a fair comparison, we remark that we used the same link predictor for CQD, CQD-Hybrid, and QTO. Details and hyperameters are available in Sec. F.1. Our analysis reveals that, similarly to the other baselines, QTO performance drops when evaluated on the full-inference $(q,a)$ pairs only (diagonal), w.r.t overall results. Moreover, even when QTO is the SoTA on a certain query type, most of the time it is not so on the portion of full-inference $(q,a)$ pairs only, showing that improving performance on the partial-inference $(q,a)$ pairs not necessarily results on improvements over the full-inference ones.

### v3

@@ 6 New Benchmarks for Undistorted CQA Performance
Building new CQA benchmarks. We generate our benchmarks to comprise both full-inference and partial-inference QA pairs, but we modify the algorithm of (Ren et al., 2020) to ensure that no training links are present in the reasoning trees of full-inference pairs, and for each query type, we build the benchmark such that the full spectrum of hardness appears with the same frequency in it. For example, for 3p queries, we sample 30,000 QA pairs, 10,000 for 3p that can be reduced to 1p, 10,000 for those reduceable to 2p, and 10,000 for full-inference QA pairs (3p QA pairs non-reduceable to any other type). Furthermore, we sample union queries 2u, 2u1p such that non-existing link are not present in their reasoning tree. Finally, to raise the bar of “complexity” in CQA, we introduce two query types that, in their full-inference versions, are harder than simpler types by design. Specifically, we design “4p” queries for a sequential path made of four links and “4i”, which represents the intersection of the target entity sets defined over four links. For their logical formulation, see App. E.
@@
How do SoTA methods fare on our new benchmarks? For $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$ we re-used the pre-trained models used for the experiments in Sec. 4, while for the newly created $\mathsf{ICEWS18}$+$\mathsf{H}$ we trained GNN-QE, CLMPT, ConE, and the link predictor used in CQD, CQD-Hybrid and QTO from scratch. As this is not necessary for ULTRAQ, being a zero-shot neural link prediction applicable to any KG, we re-used the checkpoint provided by the authors. Hyperparameters of all models are in Sec. F.2.
@@
Table 5 shows the results of the selected baselines on the new benchmarks, for all query types in Figures 2 and B.1. For stratified results on $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$ see Tables D.3, D.4 and D.5, respectively. For the stratified analysis on queries including negation see Table D.6. The surprising result from Table 5 is that no model obtains SoTA performance on most query type, which was not the case for the old benchmarks, where QTO obtained SoTA performance for most query types (see Tables A.5, A.4 and A.2). This is further evidence that current SoTA methods mainly focused on improving the performance on the easiest QA pairs and likely overfit the benchmarks. Instead, with our new benchmarks, this will not be possible in the same way, as we compile the benchmark with the same amount of partial- and full-inference QA pairs per type.
@@
Furthermore, we highlight how full-inference 2u queries have similar performance with 1p, similar to how full-inference 2u1p are expected to be as hard as 2p, as discussed in Sec. 3. Finally, we remark that $\mathsf{ICEWS18}$+$\mathsf{H}$ is much harder than $\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{FB15k237}$+$\mathsf{H}$, across all query types, even 1p, thus raising the bar for neural link predictors as well. This confirms our hypothesis that the sampling method of the KG splits plays a big role in determining the hardness of the benchmark. For additional analysis on the new benchmarks, comprising the influence of the number of existing entities bounding to existential quantified variables, the most frequent relation name and anchor entity per query type, and statistical tests, see App. D.
@@
Table 5: There is no clear SoTA method for the new benchmarks. As shown in Figure 4, the MRR on the new benchmarks is significantly lower than the old ones. For example, for 3i queries on $\mathsf{FB15k237}$+$\mathsf{H}$, QTO has an MRR of 10.1, while for $\mathsf{FB15k237}$ it was 54.6. For the stratified analysis see Tables D.3, D.4, D.5 and D.6.
Model 1p 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p 4p 4i 2in 3in 2pi1pn 2nu1p 2in1p
$\mathsf{FB15k237}$+$\mathsf{H}$ GNN-QE 42.8 5.2 4.0 6.0 8.8 5.6 9.9 32.5 10.0 4.3 20.0 6.8 6.5 3.7 5.0 3.3
ULTRAQ 40.6 4.5 3.5 5.2 7.2 5.3 10.1 29.4 8.3 3.8 16.4 5.3 5.5 2.6 3.7 2.2
CQD 46.7 4.4 2.4 11.3 12.8 6.0 11.5 40.1 10.6 1.1 23.8 3.3 2.6 0.6 4.9 1.2
CQD-Hybrid 46.7 4.8 3.1 6.0 8.6 5.5 12.9 42.2 12.0 2.4 17.4 4.7 1.6 1.0 3.2 1.3
ConE 41.8 4.6 3.9 9.1 10.3 3.8 7.9 22.8 6.0 3.5 20.3 5.1 4.9 2.9 3.3 3.6
QTO 46.7 4.9 3.7 8.7 10.1 6.1 13.5 30.6 11.2 3.9 20.2 10.6 3.1 2.0 5.3 1.5
CLMPT 45.3 5.3 4.7 10.2 12.2 5.6 14.9 33.6 14.2 4.5 24.0 6.8 2.3 1.6 4.8 2.5
$\mathsf{NELL995}$+$\mathsf{H}$ GNN-QE 53.6 8.0 6.0 10.7 13.3 16.0 13.5 47.5 9.8 4.7 19.4 5.5 6.4 5.8 3.3 4.4
ULTRAQ 38.9 6.1 4.1 7.9 10.2 15.8 9.3 28.1 9.5 4.2 15.6 4.5 5.9 4.3 2.7 3.6
CQD 60.4 9.6 4.2 18.5 19.6 18.9 22.6 46.3 18.5 2.0 24.8 4.2 1.5 1.5 4.9 2.6
CQD-Hybrid 60.4 9.0 6.1 12.1 14.4 17.4 21.2 46.4 19.3 3.5 20.4 5.1 1.2 1.4 4.3 2.4
ConE 53.1 7.9 6.7 21.8 23.6 14.9 11.8 39.9 8.8 5.2 27.6 4.6 6.0 3.7 2.7 6.4
QTO 60.3 9.8 8.0 14.6 15.8 17.6 21.1 49.1 18.9 7.0 20.9 10.2 2.3 3.1 8.4 2.4
CLMPT 58.1 10.1 7.8 22.7 25.0 17.2 24.4 50.0 22.0 7.2 29.1 6.5 2.4 4.1 2.3 4.5
$\mathsf{ICEWS18}$+$\mathsf{H}$ GNN-QE 12.2 0.9 0.5 16.1 26.5 19.1 3.5 7.6 1.1 0.4 34.0 4.5 6.9 0.9 3.5 0.8
ULTRAQ 6.3 1.2 1.2 7.0 11.7 8.8 1.3 3.3 1.2 0.8 15.9 2.3 4.8 1.2 2.2 1.6
CQD 16.6 2.5 1.5 13.0 19.5 17.1 6.7 6.8 5.9 1.1 24.0 1.5 2.9 0.2 2.7 0.9
CQD-Hybrid 16.6 2.6 1.5 15.0 25.5 17.5 5.8 6.8 5.6 0.9 33.2 1.7 4.0 0.3 2.2 1.1
ConE 3.5 0.9 0.9 1.2 0.5 1.2 1.6 1.1 0.9 0.6 0.3 1.7 2.9 1.1 0.9 1.3
QTO 16.6 2.6 1.4 15.7 25.0 18.4 6.2 6.7 4.9 1.1 31.5 4.9 8.7 1.2 3.0 0.9
CLMPT 4.7 0.8 0.1 12.0 23.0 9.7 2.1 2.7 2.2 0.1 31.0 1.2 2.1 0.1 1.0 0.2
@@
Takeaway 3. Our benchmarks are built with the same amount of partial-inference and full-inference QA pairs that depend on the query type, allowing to measure CQA performance while not distorting the aggregated performance across all QA pairs. Additionally, $\mathsf{ICEWS18}$+$\mathsf{H}$ highlights that more realistic sampling strategies are challenging for current SoTA.
@@ 7 Conclusion
In this paper, we revisit CQA on KGs and reveal that the “good” performance of SoTA approaches predominantly comes from answers that can be reduced to easier types (Table 2), the vast majority of which boiling down to single link prediction (Table 1). We perform such an analysis by dissecting the arbitrary assumptions that were used to create these datasets and by highlighting how current neural and hybrid solvers can exploit (different) forms of triple memorization to make complex queries much easier. We confirm this by reporting their performance in a stratified analysis and by proposing our hybrid solver, CQD-Hybrid, which, while being a simple extension of an old method like CQD, can be very competitive against other SoTA models. This reveals that $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ are not suitable to precisely assess the capability of CQA methods to answer complex queries and that most of their performance of CQA on $\mathsf{FB15k237}$ and $\mathsf{NELL995}$ can be explained by their memorization ability.
We then created a set of new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$$\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{ICEWS18}$+$\mathsf{H}$ that comprise the same amount of full-inference and partial-inference QA pairs. The MRR performance of existing CQA greatly drops on our new benchmarks. We consider our new benchmarks to be a stepping stone towards benchmarking truly complex reasoning with ML models and advise all future works on CQA to use them. However, both the old and our new benchmarks only consider queries with bindings of single target variables. While this reflects an important class of queries on real-world KGs, many real-world queries require bindings to multiple target variables (i.e. answer tuples). We plan to extend our current study of the “real hardness” of CQA benchmarks to other popular settings in neural query answering, such as inductive scenarios (Galkin et al., 2022; Arun et al., 2025) where some entities and/or relations are unseen during the test stage. Moreover, our work could also be extended to queries requiring bindings to multiple target variables (Yin et al., 2023), and to queries having a DAG structure (He et al., 2025).
@@ Acknowledgements
AV was supported by the “UNREAL: Unified Reasoning Layer for Trustworthy ML” project (EP/Y023838/1) selected by the ERC and funded by UKRI EPSRC. PM was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP. CG was funded by the CHIPS Joint Undertaking (JU) under grant agreement No. 101140087 (SMARTY), and by German Federal Ministry of Education and Research (BMBF) under the sub-project with the funding number 16MEE0444. CG acknowledge compute time on HoreKa HPC (NHR@KIT), funded by the BMBF and Baden-Württemberg’s MWK through the NHR program, with additional support from the DFG. BX, DH were supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1574 – 471687386.
@@ A.1 Influence of existing links on queries involving negation
In Table A.1 we show an analysis for queries involving negation. For such analysis, we split the reasoning tree, defined in Sec. 3, into two subparts, namely positive reasoning tree, composed by the non-negated triples, and the negative reasoning tree, only composed by the negated triples. In particular, in the same spirit of Table 1, by only looking at the positive reasoning tree, we report the percentage of QA pairs that can be reduced to easier types (partial-inference) and the one that cannot be reduced to a simpler type (full-inference). Furthermore, we argue that also the negative reasoning tree contains training triples, but how this propagates to performance is less clear than the positive part, as each method treats negation differently. Table A.1 shows that in both $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, queries can be reduced to easier types, revealing that our analysis also extends to negated queries. The only exceptions are 2in, 2nu1p where the positive reasoning tree consists of a single link, see App. B. Consequently, no reduction to partial inference is possible (see Sec. 3).
@@
CLMPT 7.9 6.1 3.0 11.6 7.8 4.2
@@ A.2 Additional comparisons
Table A.5 and Table A.4 show the full results of $\mathsf{FB15k237}$ and $\mathsf{NELL995}$, including the results comparing all methods, all query types, and the subtypes they are reduced to.
Moreover, A.3 show the full results of the union queries, including the old overall results, which include QA pairs having non-existing links in their reasoning tree (see Figure 3), and the “new overall”, in which those QA pairs are filtered out. In particular, for “2u1p” queries, we observe that for $\mathsf{FB15k237}$, the full-inference queries have a higher MRR than the overall, while for $\mathsf{NELL995}$, we find the MRR results on full-inference QA pairs to be lower than the new overall results as expected. We attribute the results on $\mathsf{FB15k237}$ to the presence of bias in the QA pairs due to their scarcity in the benchmarks. We argue that for the same reason, 2u1p performances are not comparable to 2p. However, we will show that for the full-inference QA pairs in the new benchmark, our claim is confirmed (see Tables D.3, D.4 and D.5).
@@
Table A.3: Full inference 2u QA pairs show higher performance than non-filtered, while full-inference 2u1p pairs show lower or comparable performance. Results on union queries for existing benchmarks and comparison between old and new overalls. ‘-’ refers to reductions that are not possible, while ‘/’ to reductions that are possible but that are not present in the data.
@@
Table A.4: The best model on all available $\mathsf{NELL995}$ queries (column “overall”) and the best on the full inference queries (diagonal) is different for 6/7 query types, excluding 1p and 2u, being already of type full-inference. Comparison of MRR scores for SoTA methods for the different subtypes $\mathsf{NELL995}$ queries can be reduced to. Best in bold.
@@
CLMPT 17.7 17.4 - - - - - - 37.5 10.2
Table A.5: The best model on all available $\mathsf{FB15k237}$ queries (column “overall”) and the best on the full inference queries (diagonal) is different for 5/7 query types, excluding 1p and 2u, those (q,a) pairs being all of type full-inference. Comparison of MRR scores for SoTA methods for the different subtypes $\mathsf{FB15k237}$ queries can be reduced to. Best in bold.
@@
The stratified MRR performance of QTO (Bai et al., 2023) for the old benchmarks have been included in Tables A.5 and A.4. For a fair comparison, we remark that we used the same link predictor for CQD, CQD-Hybrid, and QTO. Details and hyperameters are available in Sec. F.1. Our analysis reveals that, similarly to the other baselines, QTO performance drops when evaluated on the full-inference QA pairs only (diagonal), w.r.t overall results. Moreover, even when QTO is the SoTA on a certain query type, most of the time it is not so on the portion of full-inference QA pairs only, showing that improving performance on the partial-inference QA pairs not necessarily results on improvements over the full-inference ones.